efer to the following table to answer the questions that follow:CPI 80 Year 1999 2000 2001 2002 2003 2004 105 112 108 117 As presented in the table, the approximate rate of inflation (or deAs presented in the table, the approximate rate of inflation (or deflation) from 2000-2001 was (rounded to the nearest percent): A. 7%. B. 21%. C. 17%. D. 105%. E. 87%.
Answer #1

Ans) the correct option is 21%

Inflation from 2000 - 2001 = 100 * ( 105 - 87) / 87 = 20.69% or 21%